OpenStudy (anonymous):

Find Sn for this geometric series. f(1) = 1/2, r= -5, n= 6

OpenStudy (campbell_st):

well the formula for the sum of a geometric series is \[S_{n} = \frac{a(1 - r^n)}{1 - r}\] just substitute and then calculate.

OpenStudy (anonymous):

The answer in the back of my book says its -1302 :(

OpenStudy (campbell_st):

well what did you get...? so the 1st term is 1/2 so you have \[S_{6} = \frac{1/2(1 - (-5)^6)}{1 - (-5)}\] well if you calculate correctly, you'll get the answer in th back of the book,,